# MoDGS: Dynamic Gaussian Splatting from Causually-captured Monocular Videos

### v1

@@ 3.1. Gaussians and Deformation Fields
We define a set of Gaussians in the canonical space, we follow the original 3D GS (Kerbl et al., 2023) to define a 3D location, a scale vector, a rotation, and a color with spherical harmonics. Note this canonical space does not explictly correspond to any timestep but is just a virtual space that contains the canonical locations of all Gaussians.
Deformation fields
@@
The deformation field $\mathcal{T}_{t}$ used in MoDGS follows the design of Omnimotion (Wang et al., 2023a) and Cadex (Lei and Daniilidis, 2022) which is an invertible MLP network (Dinh et al., 2016). This is an invertible MLP means that both $\mathcal{T}_{t}$ and $\mathcal{T}_{t}^{-1}$ can be directly computed from the MLP network. All $\mathcal{T}_{t}$ at different timesteps $t$ share the same MLP network and the time $t$ is normalized to $[0,1]$ as input to the MLP network. We use this deformation field to deform the 3D locations of Gaussians in the canonical space but do not change the colors, orientations, scales, and rotations.
Render with MoDGS
@@
Original 3D Gaussian splatting (Kerbl et al., 2023) relies on the sparse points from Structure-from-Motion (SfM) to initialize all the locations of Gaussians. When we only have a casually captured monocular video, it is difficult to get an initial set of sparse points for initialization from SfM. Though it is possible to initialize all the Gaussians from the points of the estimated single-view depth of the first frame, we show that this leads to suboptimal results. At the same time, we need to initialize not only the Gaussians but also the deformation field. Thus, in the following, we propose a 3D-aware initialization scheme for MoDGS.
Initialize depth scales
@@
Since the estimated depth maps on different timesteps would have different scales, we first estimate a coarse scale for every frame to unify the scales. We consider the scale of the first frame as 1.0 and normalize all other depth maps to the first frame. Since the input video is captured by a casual camera almost static without extremely large movements, we segment the pseudo-static regions on every timestep by thresholding on 2D flow $F_{t\to t+1}$. Then, on these pseudo-static regions, we reproject the depth values at a specific timestep to the first frame and minimize the difference between the projected depth and the depth of the first frame, which enables us to solve for a scale for every frame. We rectify all depth maps with the computed scales. In the following, we reuse $D_{t}$ to denote the rectified depth maps by default.
Initialization of the deformation field
@@
As shown in Fig. 3 (left), given two depth maps $D_{t_{i}}$ and $D_{t_{j}}$ along with the 2D flow $F_{t_{i}\to t_{j}}$, we lift them to a 3D flow $F_{t_{i}\to t_{j}}^{3D}$. This is achieved by first converting the depth maps into 3D points in the 3D space. Then, the estimated 2D flow $F_{t_{i}\to t_{j}}$ actually associate two sets of 3D points, which results in a 3D flow $F_{t_{i}\to t_{j}}^{3D}$. After getting this 3D flow, we then train our deformation field $\mathcal{T}$ with this 3D flow. Specifically, for a pixel in $I_{t_{i}}$ whose corresponding 3D point is $x_{t_{i}}$, we query $F^{3D}_{t_{i}\to t_{j}}$ to find its target point $x_{t_{j}}$ in the $t_{j}$ timestep. Then, we minimize the difference by
(1) $\ell_{\text{init}}=\sum\|\mathcal{T}_{t_{j}}\circ\mathcal{T}^{-1}_{t_{i}}(x_{t_{i}})-x_{t_{j}}\|^{2}.$
@@
We train the MLP networks in $\mathcal{T}$ for a fixed number of steps to initialize the deformation field.
Initialization of Gaussians
@@
After getting the initialized deformation field, we will initialize a set of 3D Gaussians in the canonical space as shown in Fig. 3 (right). We achieve this by first converting all the depth maps to get 3D points. Then, these 3D points are deformed backward to the canonical 3D space. This means that we transform all the depth points of all timesteps to the canonical space, which results in a large amount of points. We then evenly downsample these points with a predefined interval to reduce the point number and we initialize all our Gaussians with the locations of these downsampled 3D points in the canonical space.
3.3. Ordinal Depth Loss
Pearson correlation loss
@@ 3.3. Ordinal Depth Loss
(2) $\operatorname{Corr}\left(\hat{D}_{t},D_{t}\right)=\frac{\operatorname{Cov}\left(\hat{D}_{t},D_{t}\right)}{\sqrt{\operatorname{Var}\left(\hat{D}_{t}\right)\operatorname{Var}\left(D_{t}\right)}},$
where $\operatorname{Cov}\left(\cdot,\cdot\right)$ and $\operatorname{Var}\left(\cdot,\cdot\right)$ means the covariance and variance respectively, $D_{t}$ and $\hat{D}_{t}$ are the estimated depth and the rendered depth respectively. Since the estimated single-view depth has ambiguity in scale, the Pearson correlation loss avoids the negative effects of the scale ambiguity. Note that in (Li et al., 2021; Liu et al., 2023a), the loss is called normalized depth loss, which is equivalent to Pearson correlation here as shown in the supplementary material.
Problem in Pearson correlation loss
However, we find that this Pearson correlation depth loss is still suboptimal. As shown in Fig. 4, the estimated depth maps at two different timesteps are still not consistent with each other after normalization. Making two depth maps consistent after normalization actually requires these two depth maps to be related by a linear transformation, i.e. $D_{t+1}=aD_{t}+b$ with $a$ and $b$ two constants. However, the single-view depth estimation method is not accurate enough to guarantee the linear relationship between two estimated depth maps at different timesteps. In this case, the Pearson correlation loss still brings inconsistent supervision for training the 3D representation.
Ordinal depth loss
@@
(3) ${\mathcal{R}}(D_{t}(u_{1}),D_{t}(u_{2}))=\begin{cases}+1,&D_{t}(u_{1})>D_{t}(u_{2})\\ -1,&D_{t}(u_{1})<D_{t}(u_{2})\\ 0,&D_{t}(u_{1})=D_{t}(u_{2})\end{cases},$
@@
(4) $\ell_{\text{ordinal}}=\|\text{tanh}\left(\alpha(\hat{D}_{t}(u_{1})-\hat{D}_{t}(u_{2}))\right)-\mathcal{R}\left(D_{t}(u_{1}),D_{t}(u_{2})\right)\|,$
@@
where $\text{tanh}(x)=\frac{{e^{x}-e^{-x}}}{{e^{x}+e^{-x}}}$, $\hat{D}_{t}$ means the rendered depth map at time step $t$, $\hat{D}_{t}(u)$ is the depth value of this rendered depth map on the pixel $u$, $\alpha$ is a predefined constant. Eq. 4 means we transform the depth difference between $\hat{D}_{t}(u_{1})$ and $\hat{D}_{t}(u_{2})$ to 1 or -1 by tanh function. Then, we force the depth order of the rendered depth map $\hat{D}_{t}$ to be consistent with the order in the predicted depth map $D_{t}$. In the implementation, we randomly sample 100k pairs $(u_{1},u_{2})$ to compute the ordinal depth loss.
Figure 4. We show the estimated single-view depth maps at two different timesteps $D_{t_{i}}$ and $D_{t_{j}}$ after normalization to the same scale. Since the single-view depth estimator is not accurate enough, the depth maps are not linear related so the scale normalization does not perfectly align them. However, the order of depth values on three corresponding pixels is stable for these two depth maps, which motivates us to propose an ordinal depth loss for supervision.
3.4. Training of MoDGS
@@
After initializing the Gaussians and the deformation fields, we use MoDGS to render at a specific timestep and compute the rendering loss $\ell_{\text{render}}$ and the ordinal depth loss $\ell_{\text{ordinal}}$. So the total training loss for MoDGS is
(5) $\ell=\lambda_{\text{ordinal}}\ell_{\text{ordinal}}+\lambda_{\text{render}}\ell_{\text{render}}.$
4. Experiments
@@
Figure 5. Qualitative comparison on the novel-view renderings of the DyNeRF (Li et al., 2022) and Nvidia (Yoon et al., 2020) datasets. We compare MoDGS with SC-GS (Huang et al., 2024), Deformable-GS (Yang et al., 2023a), and Hexplane (Cao and Johnson, 2023).
4.1. Implementation Details
We implement our MoDGS with PyTorch. To initialize the deformation field, we train it with 30k steps as stated in Sec. 3.2. Subsequently, we jointly train the 3D Gaussians and the deformation field with the rendering loss and the ordinal depth loss for another 30k steps. The downsampling voxel size for Gaussian initialization is 0.004. We adopt an Adam optimizer for optimization. The learning rate for 3D Gaussians exactly follows the official implementation of 3D GS (Kerbl et al., 2023), while the learning rate of the deformation network undergoes exponential decay from 1e-3 to 1e-4 in initialization and from 1e-4 to 1e-6 in the subsequent optimization. We set $\alpha=100$ for $\ell_{\text{ordinal}}$. The weight of our depth order loss is 0.1. The whole training takes around 6 hours to converge (3 hours for the initialization and 3 hours for the subsequent optimization) on an NVIDIA RTX A6000 GPU, which uses about 14G memory. The rendering speed of MoDGS is about 70 FPS.
@@ 4.1. Implementation Details
Figure 6. Qualitative comparison of DVS quality on the MCV dataset. We compare MoDGS with SC-GS (Huang et al., 2024), Deformable-GS (Yang et al., 2023a), Hexplane (Cao and Johnson, 2023), and RoDynRF (Liu et al., 2023a).
4.2. Evaluation Protocols
Datasets
@@ 4.2. Evaluation Protocols
We conducted experiments on three datasets to demonstrate the effectiveness of our method. The first dataset is the DyNeRF (Li et al., 2022) dataset which consists of 6 scenes. On each scene, we have 18-20 synchronized cameras capturing 10-30 second videos. In these videos, there is mainly a man working on a desktop, like cutting beef or dumping water. We use 5 scenes for the evaluation of the DyNeRF dataset. We use camera0 for training and evaluate the results on camera5 and camera6. The second dataset is the Nvidia (Yoon et al., 2020) dataset which contains more diverse dynamic subjects like jumping, playing with balloons, and opening an umbrella. The Nvidia dataset contains 8 scenes, which also has 12 synchronized cameras. We use 3 scenes for quantitative evaluation. We train all methods on camera4 and evaluate with camera3 and camera5. Other than these widely-used benchmarks, we also collect 6 online videos to construct an in-the-wild dataset, called the Monocular Casual Video (MCV) Dataset, to demonstrate our method can generalize to in-the-wild casual videos. The MCV dataset contains diverse subjects like skating, a dog eating food, YOGA, etc. The MCV dataset only contains a single video for each scene, so we cannot evaluate the quantitative results but only report the qualitative results on this dataset.
Evaluation setting
@@
Previous DVS methods (Cao and Johnson, 2023; Yang et al., 2023a; Gao et al., 2021) all use different cameras to train the dynamic NeRF or Gaussian Splatting. Even though they only use one camera at a specific timestep, they use different cameras at different timesteps so that a pseudo multiview video can be constructed to learn the 3D structures of the scene. Since our target is to conduct novel-view synthesis on the casually captured images, we do not adopt such ”teleporting camera motions” to construct training videos but just adopt one static camera to record training videos. Then, we render the images from the viewpoints of another camera for evaluation.
Metrics
@@
Nvidia Dataset
Ours SC-GS D-GS Hexplane
@@
balloon2-2 20.47 0.5275 0.2408 18.28 0.3893 0.3671 18.57 0.3941 0.3768 18.42 0.3159 0.4881
skatting2 25.64 0.7996 0.1518 23.07 0.7186 0.2002 24.47 0.7582 0.2073 21.39 0.6403 0.3983
truck-2 23.69 0.7455 0.1551 21.48 0.6374 0.2078 21.38 0.6321 0.2142 20.85 0.5687 0.3653
Avg. 23.27 0.6908 0.1826 20.94 0.5817 0.2584 21.47 0.5948 0.2661 20.22 0.5083 0.4172
[…]
DyNeRF Dataset
Ours SC-GS D-GS HexPlane
[…]
cut_beef 23.98 0.8221 0.1438 20.69 0.7414 0.2625 22.20 0.7808 0.1931 16.76 0.5382 0.5054
sear_steak 23.53 0.8126 0.1642 21.23 0.7870 0.2188 23.56 0.8101 0.1773 16.89 0.5897 0.5049
coffee_martini 21.37 0.7962 0.1473 19.02 0.7124 0.2151 19.23 0.7013 0.2270 13.26 0.4049 0.5835
cook_spinach 22.40 0.7823 0.1728 16.70 0.7377 0.2117 17.20 0.7195 0.2329 16.95 0.7286 0.2223
flame_steak 23.23 0.8083 0.1592 17.31 0.7532 0.2527 16.62 0.7523 0.2559 16.97 0.7528 0.2543
Avg. 22.90 0.8043 0.1575 18.99 0.7463 0.2322 19.76 0.7528 0.2172 16.17 0.6028 0.4141
Baseline methods
@@
We compare MoDGS with 4 baseline methods to demonstrate the superior ability of MoDGS to synthesize novel-view images with casually captured monocular videos. These methods can be categorized into two classes. The first is the NeRF-based methods including HexPlane (Cao and Johnson, 2023) and RoDynRF (Liu et al., 2023a). HexPlane represents the scene with six feature planes in both 3D space and time-space. We find that HexPlane does not produce reasonable results if only a monocular video from a single camera is given as input. Thus, other than the input monocular video, we use another video from a different viewpoint to train HexPlane for the DVS task. RoDynRF is a SoTA NeRF-based DVS method that also adopts single-view depth estimation as supervision for the 3D dynamic representations. We train it with the same single-view depth estimator GeoWizard (Fu et al., 2024) as ours. The second class is the Gaussian Splatting-based DVS methods including Deformable-GS (Yang et al., 2023a) and SC-GS (Huang et al., 2024). Deformable-GS also associates a deformation field with a set of canonical Gaussians for DVS. SC-GS learns a set of keypoints and uses the deformation of these keypoints to blend the deformation of arbitrary 3D points.
4.3. Comparison with Baselines

### v2

@@ 3.1 Gaussians and Deformation Fields
We define a set of Gaussians in the canonical space, we follow the original 3D GS (Kerbl et al., 2023) to define a 3D location, a scale vector, a rotation, and a color with spherical harmonics. Note this canonical space does not explictly correspond to any timestamp but is just a virtual space that contains the canonical locations of all Gaussians.
Deformation fields.
@@
The deformation field $\mathcal{T}_{t}$ used in MoDGS follows the design of Omnimotion (Wang et al., 2023c) and CaDeX (Lei & Daniilidis, 2022) which is an invertible MLP network (Dinh et al., 2016). This is an invertible MLP means that both $\mathcal{T}_{t}$ and $\mathcal{T}_{t}^{-1}$ can be directly computed from the MLP network. All $\mathcal{T}_{t}$ at different timestamps $t$ share the same MLP network and the time $t$ is normalized to $[0,1]$ as input to the MLP network.
Render with MoDGS.
@@
Original 3D Gaussian splatting (Kerbl et al., 2023) relies on the sparse points from Structure-from-Motion (SfM) to initialize all the locations of Gaussians. When we only have a casually captured monocular video, it is difficult to get an initial set of sparse points for initialization from SfM. Though it is possible to initialize all the Gaussians from the points of the estimated single-view depth of the first frame, we show that this leads to suboptimal results. At the same time, we need to initialize not only the Gaussians but also the deformation field. Thus, we propose a 3D-aware initialization scheme for MoDGS.
Initialization of depth scales.
Since the estimated depth maps on different timestamps would have different scales, we first estimate a coarse scale for every frame to unify the scales. We achieve this by first segmenting out the static regions on the video and then computing the scale with a least square fitting (Chung et al., 2023b). The static regions can be determined by either thresholding on the 2D flow (Teed & Deng, 2020) or segmenting with a segmentation method like SAM2 (Ravi et al., 2024).
@@
Then, on these static regions, we reproject the depth values at a specific timestamp to the first frame and minimize the difference between the projected depth and the depth of the first frame, which enables us to solve for a scale for every frame. We rectify all depth maps with the computed scales. In the following, we reuse $D_{t}$ to denote the rectified depth maps by default.
Initialization of the deformation field.
@@
As shown in Fig. 3 (left), given two depth maps $D_{t_{i}}$ and $D_{t_{j}}$ along with the 2D flow $F_{t_{i}\to t_{j}}$, we lift them to a 3D flow $F_{t_{i}\to t_{j}}^{3D}$. This is achieved by first converting the depth maps into 3D points in the 3D space. Then, the estimated 2D flow $F_{t_{i}\to t_{j}}$ actually associate two sets of 3D points, which results in a 3D flow $F_{t_{i}\to t_{j}}^{3D}$. After getting this 3D flow, we then train our deformation field $\mathcal{T}$ with this 3D flow. Specifically, for a pixel in $I_{t_{i}}$ whose corresponding 3D point is $x_{t_{i}}$, we query $F^{3D}_{t_{i}\to t_{j}}$ to find its target point $x_{t_{j}}$ in the $t_{j}$ timestamp. Then, we minimize the difference by
$\ell_{\text{init}}=\sum\|\mathcal{T}_{t_{j}}\circ\mathcal{T}^{-1}_{t_{i}}(x_{t_{i}})-x_{t_{j}}\|^{2}.$ (1)
@@
We train the MLP in $\mathcal{T}$ for a fixed number of steps to initialize the deformation field.
Initialization of Gaussians.
@@
After getting the initialized deformation field, we will initialize a set of 3D Gaussians in the canonical space as shown in Fig. 3 (right). We achieve this by first converting all the depth maps to get 3D points. Then, these 3D points are deformed backward to the canonical 3D space. This means that we transform all the depth points of all timestamps to the canonical space, which results in a large amount of points. We then evenly downsample these points with a predefined voxel size to reduce the point number and we initialize all our Gaussians with the locations of these downsampled 3D points in the canonical space.
3.3 Ordinal Depth Loss
Pearson correlation loss.
@@ 3.3 Ordinal Depth Loss
$\operatorname{Corr}\left(\hat{D}_{t},D_{t}\right)=\frac{\operatorname{Cov}\left(\hat{D}_{t},D_{t}\right)}{\sqrt{\operatorname{Var}\left(\hat{D}_{t}\right)\operatorname{Var}\left(D_{t}\right)}}$,
where $\operatorname{Cov}\left(\cdot,\cdot\right)$ and $\operatorname{Var}\left(\cdot,\cdot\right)$ means the covariance and variance respectively, $D_{t}$ and $\hat{D}_{t}$ are the estimated depth and the rendered depth respectively. Since the estimated single-view depth has ambiguity in scale, the Pearson correlation loss avoids the negative effects of the scale ambiguity. Note that in  Li et al. (2021) and  Liu et al. (2023a), the loss is called normalized depth loss, which is equivalent to Pearson correlation here as shown in the supplementary material.
Figure 4: We show the estimated single-view depth maps at two different timestamps $D_{t_{i}}$ and $D_{t_{j}}$ after normalization to the same scale. Since the single-view depth estimator is not accurate enough, the depth maps are not linear related so the scale normalization does not perfectly align them. However, the order of depth values on three corresponding pixels is stable for these two depth maps, which motivates us to propose an ordinal depth loss for supervision.
Limitations of Pearson correlation loss.
However, we find that this Pearson correlation depth loss is still suboptimal. As shown in Fig. 4, the estimated depth maps at two different timestamps are still not consistent with each other after normalization. Making two depth maps consistent after normalization actually requires these two depth maps to be related by a linear transformation, i.e. $D_{t+1}=aD_{t}+b$ with $a$ and $b$ two constants. However, the single-view depth estimation method is not accurate enough to guarantee the linear relationship between two estimated depth maps at different timestamps. In this case, the Pearson correlation loss still brings inconsistent supervision.
Ordinal depth loss.
@@
${\mathcal{R}}(D_{t}(u_{1}),D_{t}(u_{2}))=\begin{cases}+1,&D_{t}(u_{1})>D_{t}(u_{2})\\ -1,&D_{t}(u_{1})<D_{t}(u_{2})\\ \end{cases},$ (2)
@@
$\ell_{\text{ordinal}}=\|\text{tanh}\left(\alpha(\hat{D}_{t}(u_{1})-\hat{D}_{t}(u_{2}))\right)-\mathcal{R}\left(D_{t}(u_{1}),D_{t}(u_{2})\right)\|,$ (3)
@@
where $\text{tanh}(x)=\frac{{e^{x}-e^{-x}}}{{e^{x}+e^{-x}}}$, $\hat{D}_{t}$ means the rendered depth map at timestamp $t$, $\hat{D}_{t}(u)$ is the depth value of this rendered depth map on the pixel $u$, $\alpha$ is a predefined constant. Eq. 3 means we transform the depth difference between $\hat{D}_{t}(u_{1})$ and $\hat{D}_{t}(u_{2})$ to 1 or -1 by tanh function. Then, we force the depth order of the rendered depth map $\hat{D}_{t}$ to be consistent with the order in the predicted depth map $D_{t}$. In the implementation, we randomly sample 100k pairs $(u_{1},u_{2})$ to compute the ordinal depth loss.
3.4 Training of MoDGS
@@
After initializing the Gaussians and the deformation fields, we use MoDGS to render at a specific timestamp and compute the rendering loss $\ell_{\text{render}}$ and the ordinal depth loss $\ell_{\text{ordinal}}$. So the total training loss for MoDGS is
$\ell=\lambda_{\text{ordinal}}\ell_{\text{ordinal}}+\lambda_{\text{render}}\ell_{\text{render}}.$ (4)
4 Experiments
@@
Figure 5: Qualitative comparison on the novel-view renderings of the DyNeRF (Li et al., 2022) and Nvidia (Yoon et al., 2020) datasets. We compare MoDGS with SC-GS (Huang et al., 2024), Deformable-GS (Yang et al., 2023b), and HexPlane (Cao & Johnson, 2023).
@@
Figure 6: Qualitative comparison of DVS quality on the MCV dataset. We compare MoDGS with SC-GS (Huang et al., 2024), Deformable-GS (Yang et al., 2023b), HexPlane (Cao & Johnson, 2023), and RoDynRF (Liu et al., 2023a).
4.1 Evaluation Protocols
Datasets.
@@ 4.1 Evaluation Protocols
We conducted experiments on four datasets to demonstrate the effectiveness of our method. The first dataset is the DyNeRF (Li et al., 2022) dataset which consists of 6 scenes. On each scene, we have 18-20 synchronized cameras capturing 10-30 second videos. In these videos, there is mainly a man working on a desktop, like cutting beef or dumping water. We use camera0 for training and evaluate the results on camera5 and camera6. The second dataset is the Nvidia (Yoon et al., 2020) dataset which contains more diverse dynamic subjects like jumping, playing with balloons, and opening an umbrella. The Nvidia dataset contains 8 scenes, which also has 12 synchronized cameras. We train all methods on camera4 and evaluate with camera3 and camera5. Besides, we also collect 6 online videos to construct an in-the-wild dataset, called the Monocular Casual Video (MCV) Dataset, to demonstrate our method can generalize to in-the-wild casual videos. The MCV dataset contains diverse subjects like skating, a dog eating food, YOGA, etc. The MCV dataset only contains a single video for each scene, so we cannot evaluate the quantitative results but only report the qualitative results on this dataset. We also present results of the Davis dataset (Pont-Tuset et al., 2017) in Sec. A.6 of the appendix.
Evaluation setting.
@@
Previous DVS methods (Cao & Johnson, 2023; Yang et al., 2023b; Gao et al., 2021) all use different cameras to train the dynamic NeRF or Gaussian Splatting. Even though they only use one camera at a specific timestamp, they use different cameras at different timestamps so that a pseudo multiview video can be constructed to learn the 3D structures of the scene. Since our target is to conduct novel-view synthesis on the casually captured images, we do not adopt such ”teleporting camera motions” to construct training videos but just adopt one static camera to record training videos. Then, we render the images from the viewpoints of another camera for evaluation.
Metrics.
@@
DyNeRF Nvidia
@@
HexPlane 15.33 0.5593 0.4514 17.17 0.3675 0.4756
SC-GS 18.77 0.7359 0.2310 17.59 0.4679 0.3348
D-GS 19.55 0.7446 0.2171 18.07 0.4650 0.3422
Ours 22.64 0.8042 0.1545 19.27 0.5235 0.2581
Baseline methods.
@@
We compare MoDGS with 4 baseline methods to demonstrate the superior ability of MoDGS to synthesize novel-view images with casually captured monocular videos. These methods can be categorized into two classes. The first is the NeRF-based methods including HexPlane (Cao & Johnson, 2023) and RoDynRF (Liu et al., 2023a). HexPlane represents the scene with six feature planes in both 3D space and time-space. We find that HexPlane does not produce reasonable results if only a monocular video from a single camera is given as input. Thus, other than the input monocular video, we use another video from a different viewpoint to train HexPlane for the DVS task. RoDynRF is a SoTA NeRF-based DVS method that also adopts single-view depth estimation as supervision for the 3D dynamic representations. We train it with the same single-view depth estimator GeoWizard (Fu et al., 2024) as ours. The second class is the Gaussian Splatting-based DVS methods including Deformable-GS (Yang et al., 2023b) and SC-GS (Huang et al., 2024). Deformable-GS also associates a deformation field with a set of canonical Gaussians for DVS. SC-GS learns a set of keypoints and uses the deformation of these keypoints to blend the deformation of arbitrary 3D points.
4.2 Comparison with Baselines